# Denoising Vision Transformers

### v1

@@ Appendix C Discussion on Limitations
Our work serves as one of the initial studies into understanding the position-based artifacts present in the features of ViT models. We explore the presence of such artifacts and propose a method to denoise such artifacts. However, the underlying reason for why such artifacts exist, and in which way, remains elusive. In particular, the severity of the artifacts depends on the algorithm that it is trained on, i.e. DINOv2 has more exaggerated artifacts while MAE has weaker artifacts. Thus, one direction of exploration is investigating the training paradigm including the supervision —i.e. local vs. global — as well as the loss-induced parameter landscape —i.e. sharp vs. smooth Hessians. Furthermore, a better architectural design—e.g. new positional embeddings—may diminish the severity of the feature artifacts. In this work, we do not explore modifying the ViT’s design; however, more study into its positional embeddings and the effect on downstream features should prove interesting. Ultimately, we believe our findings are intriguing to the community and further research is needed to better understand this fundamental problem.

### v2

@@ Appendix D Discussion on Limitations
Our work serves as one of the initial studies to understand the position-based artifacts present in the features of ViT models. We identify and propose methods to mitigate these artifacts, yet the root causes and characteristics of these artifacts are not fully understood. The severity of artifacts varies with the training algorithms; for instance, DINOv2 exhibits more pronounced artifacts compared to MAE, which shows subtler discrepancies. Thus, one direction of exploration is to investigate the training paradigm that includes supervision —i.e. local vs. global — as well as the loss-induced parameter landscape —i.e. sharp vs. smooth Hessians. Furthermore, a better architectural design—e.g. new positional embeddings—may diminish the severity of the feature artifacts. In this work, we do not explore modifying the ViT’s design; however, more study into its positional embeddings and the effect on downstream features should prove interesting. Ultimately, we believe our findings are intriguing to the community and more research is needed to better understand this fundamental problem.